The Bureau of Alcohol, Tobacco, and Firearms (BATF) has been concerned about lead levels in California wines. In a previous testing of wine specimens, lead levels ranging from 47 to 700 parts per billion were recorded. How many wine specimens should be tested if the BATF wishes to estimate the true mean lead level for California wines to within 10 parts per billion with 95% confidence? (Round your answer up to the next whole number.)
Answer #1

Solution :

Range = maximum - minimum = 700 - 47 = 653

Using Range rule of thumb

Standard deviation = Range / 4 = 653 / 4 = 163.25

Margin of error = E = 10

Z/2 = 1.96

sample size = n = [Z/2* / E] 2

n = [1.96 * 163.25 / 10]2

n = 1024

Sample size = n = 1024